# algorithmic

### v2

@@ 3.2.3 Representation-Specific Classes
While many approaches either rely on hard-coded features or learn a model entirely end-to-end, it is also possible to actively elicit new features from human feedback. Feature traces were proposed by Bobu et al. (2022) as an approach to actively learn new relevant features. This type of feedback relies on a human operator to demonstrate a behavior in which a certain feature of interest, such as the distance to a sensitive object, monotonically increases or decreases. They make it possible to extend the set of features once the current set can no longer adequately explain the human feedback supplied through another type of feedback. The defining characteristics of feature traces are that they are of descriptive (intent) and explicitly (explicitness) given in an active manner (involvement) for a single (arity) segment (granularity) on an instance-level abstraction.
Feature traces are strongly related to inter-temporal preferences (Section 3.2.1) since both types rely on changes in feature or reward values in the course of a single trajectory. Bobu et al. (2022) propose to learn from feature traces by leveraging a Bradley-Terry model to learn the feature values, similar to other approaches that use such a model to learn reward values. Similar to importance feedback, feature traces rely on another type of feedback to actually make use of the learned features and is, therefore, a purely supplementary form of feedback. For instance, Bobu et al. (2022) use intervention feedback to train a reward model on the set of features derived using feature traces.
@@
Similarity queries are a feedback type aimed at learning a representation conforming to a notion of similarity and difference in the trajectory space. That aim is closely aligned with that of feature queries, though the actual queries are more similar to comparisons. The queries consist of triples of trajectories, with one anchor and two alternatives, for which the human has to decide which pair is more similar. Responses to similarity queries are given on observed behavior (involvement) with ternary arity, descriptive intent, and explicit feedback, while the granularity and abstraction may vary. This type of feedback was first introduced by Bobu et al. (2023), who used it to learn representations for reward learning.
@@ 3.3 Initializations
Some modes of communicating reward functions are not interactive nor online and, therefore, do not directly fit within the scope of this survey (Section 1.3). Since these are often used to initialize a reward function for later refinement with some of the previously discussed interactive feedback types, they are still worth mentioning.
Initializations are most commonly given by examples of successful task completions, either in the form of terminal or goal states (Xie et al., 2018), expert demonstrations (Ibarz et al., 2018; Fu et al., 2018a; Palan et al., 2019; Lee et al., 2021b; Bıyık et al., 2022a; Abramson et al., 2022; Huang et al., 2023), labeled demonstrations (Du et al., 2023), or ranked demonstrations (Brown et al., 2019). It is even possible to infer some human preferences by the state of the environment alone, e.g., assuming the parts of the initial environment that are under the user’s control are largely set up in accordance with the user’s preferences (Shah et al., 2019; Lindner et al., 2021a). Since offline initialization is not the main focus of our survey, we do not cover these works in detail. We refer the interested reader to literature on inverse RL (Arora & Doshi, 2021) for further details on learning from demonstrations in particular.
@@ 3.4 Choice of Feedback Type
The best feedback type is not always clear and may depend on the task, the user, or the agent. It may also change over time as the agent learns more about the task. Table 4 may serve as a starting point to select a set of feedback types which may be applicable based on the possible user interaction and expertise, e.g., by the desired granularity, level of abstraction or involvement. This choice may be informed, among other features, by the task complexity or time horizon: Jain et al. (2015) compare purely passive (re-ranking based) with active (slight trajectory improvements) feedback and conclude that the former is usually preferred by humans when the action space is large while the latter is preferred for complex tasks. Similarly, Sumers et al. (2023) empirically demonstrate that language is a more effective teaching modality than demonstrations for complex concepts. The relevance of the time horizon is highlighted in the work by Sumers et al. (2022b), which suggests that instructive feedback is more efficient for short time horizons and descriptive feedback for longer horizons.
In addition to these static choices, Section 4.1.2 will discuss how to choose a feedback type adaptively. It is also possible to let the user choose the feedback type, as demonstrated by Jain et al. (2015) who let the user choose between re-ranking and improvement. Jeon et al. (2020) propose to use this choice of feedback type itself as a source of information (“meta-choice”).
@@ 3.5 Combination of Feedback Types
The two-phase setup can be used to either initialize the reward model from offline data or to learn a representation that improves later reward learning. A common approach involves using one feedback type, typically demonstrations (see Section 3.3), to initialize the reward model, subsequently fine-tuning this model with another type, such as comparisons. This method is exemplified by Ibarz et al. (2018), who combined demonstrations and pairwise comparisons. For a more detailed discussion on feedback types suitable for initialization, refer to Section 3.3. Alternatively, a representation-focused type of feedback might be employed initially (Section 3.2.3) to cultivate a superior representation, followed by the application of a primary feedback type for reward model learning.
Combining a primary feedback type with a supplementary one can be beneficial to make the most out of the available user-interactions. Supplementary feedback, while not sufficient for reward model training by itself, can often be collected cheaply or as a side-effect of other interactions, making it a valuable addition to the primary feedback type. Refer to Section 3.2.2 for a discussion on supplementary feedback types.
@@ 4 Label Collection
Figure 3: Highlighting the components of the RLHF loop discussed in this section.
In this section, we explain how preference data can be collected for training a reward model, independent of the specific type of feedback used. We start by overviewing the active learning problem posed by query selection, e.g., how queries can be generated for a given query type, and how even the type of feedback itself can be selected. Following this, we discuss issues arising in such human-computer interaction.
4.1 Active Learning
We first discuss query generation and selection for a given feedback type, then the extension of these techniques to the choice of feedback types.
4.1.1 Query Generation and Selection
As the RL agent is trained and its learned policy changes, the trajectories it generates will naturally evolve. Most work directly uses the trajectories obtained during RL training for preferences learning. Using such trajectories, candidate queries are often randomly generated. However, for pairwise comparison, a more efficient approach is to generate queries by exploiting preference transitivity (Hwang et al., 2023). Alternatively, the agent can also generate trajectories specifically to be used for querying (not necessarily for RL training), possibly with a learned transition model to limit the sampling cost (e.g., Reddy et al. (2020); Liu et al. (2023b)). This kind of active generation of queries can possibly lead to more informative ones.
@@ 4.1.1 Query Generation and Selection
Reddy et al. (2020) Ensemble averaged KL-
@@
Hu et al. (2024) Random
In order to efficiently learn a suitable reward model, the agent must generate and select queries (Line 5 from Algorithm 1) so that it can quickly learn a good strategy using those queries. This selection is performed via a criterion, usually called acquisition function [^1] We follow the terminology used in Bayesian active learning., which allows the queries to be compared. Although most existing work in RLHF uses an acquisition function that provides some measure of uncertainty (e.g., about the learned rewards), which is arguably one of the most important factors in active learning, an efficient acquisition function (see Table 5 for works dedicated to improving query selection) may need to include various additional other aspects, such as: on-policy data, query simplicity, trajectory quality, query diversity, or query cost, which will be discussed one by one in the following. As a side note, interestingly, as highlighted by Habibian et al. (2022), the queries asked by an RL agent may also reveal its current reward learning stage.
@@
Finally, it should be mentioned that naturally approaches with theoretical guarantees usually also use uncertainty as the main criterion for label collection (e.g., Novoseller et al. (2023), see Section 7 for more details).
@@
Only focusing on uncertainty is likely insufficient or inefficient in RL because the previous methods may focus on choosing queries to identify the reward function as precisely as possible uniformly on the whole state-action space. However, it may be important to favor more on-policy trajectories to guarantee the relevance of the generated queries for the current policy. Indeed, improving reward learning in state-action regions that may never be visited with the current policy would lead to wasteful queries (Lindner et al., 2021b). One simple approach to ensure that the data is more on-policy is by simply sampling from the current policy (Cui & Niekum, 2018) or favoring more recently-generated trajectories (Eberhard et al., 2022; Hu et al., 2024).
@@
When asking many queries (in batch, in sequence), the diversity of the queries becomes especially crucial to avoid asking redundant queries. Most work (Christiano et al., 2017; Lee et al., 2021b; Verma & Metcalf, 2024) follows a very myopic approach: Queries are often selected from a usually randomly-generated set of potential queries, and sequences of queries are not really coordinated. While some work exists that specifically tackles the selection of a batch of diverse queries (Bıyık & Sadigh, 2018; Bıyık et al., 2024), the latter is rarely considered due to its computational intractability. Indeed, planning ahead a sequence of queries would amount to solving a sequential decision-making problem under uncertainty over a combinatorial action space (i.e., the set of possible queries). For diverse batch querying, previous work considered using clustering methods such as k-medoids (Bıyık & Sadigh, 2018) or more recently determinantal point processes, which define probability distributions that promote diversity (Bıyık et al., 2024).
@@
Since various different acquisition functions have been considered, some effort (Lee et al., 2021b; a) has been made to compare them. Generally speaking, uncertainty-based criteria (e.g., variance or average entropy) seem to often perform better empirically compared to random selection, a query diversity-based criterion alone or combined with an uncertainty-based criterion. In addition, Eberhard et al. (2022) empirically observe that a variance-based criterion performs better than a selection method only based on trajectory recency. Surprisingly, random selection has been shown to perform competitively in some cases (Christiano et al., 2017; Ibarz et al., 2018). Thus, a better understanding of which acquisition function should be preferred in which situation or domain is still an open question.
@@
where a higher $\beta$ corresponds to a more reliable human (see Section 5.1.2). The setting in which this assumption does not hold, i.e., the labeler’s reward functions differ (a setting already considered in inverse RL (Choi & Kim, 2012)), has also been studied recently (Siththaranjan et al., 2024; Xue et al., 2023a; Dong et al., 2024; Myers et al., 2021; Bakker et al., 2022). Interestingly, a noisy oracle may sometimes provide more information than a completely reliable oracle. For instance, in Eq. 3, the frequency of erroneous answers given by the noisy oracle is related to how much a segment is preferred to the other one (Xu et al., 2020; Chan et al., 2021). In contrast, only a binary preorder over segments can be inferred from the answers of a reliable and deterministic oracle, which may not be enough to recover the true reward function.
Another notable recent work is that by Ellis et al. (2024) who raise the issue of identifiability of the ground-truth reward function: many reward functions result in the same optimal behaviors. The authors propose a framework that enables the generation of acquisition function for various definitions of reward similarity, such as the one discussed in Section 5.3.
@@ 4.1.2 Adaptive Choice of Feedback Type
[…]
@@
Ghosal et al. (2023) find that the most informative feedback type depends on the (type-dependent) rationality of the human labeler (see Section 4.2.1). More precisely, it is shown that the most informative feedback depends on the rationality factor, e.g., while demonstrations are more informative than comparisons when the human is highly rational, comparisons should be preferred in less-rational settings. Given that this rationality might change due to factors such as fatigue or an individual labeler’s capabilities, this suggests that adaptively adjusting the feedback type during the labeling process may be worthwhile. Further study of this relationship is a promising area for future work.
@@ 4.2.1 Psychology-Aware Preference Elicitation
Preference learning methods typically assume the existence of inherent, stable preferences that can be elicited through querying. Contrary to this assumption, psychological research, such as the work by Lichtenstein & Slovic (2006), indicates that preferences are often constructed during the elicitation process and may vary with the method of elicitation or over time. This suggests that the feedback type not only affects elicitation’s effectiveness but also shapes preferences. Systematic biases, noise, and other psychological factors may influence observed choices, challenging the traditional models of human choice used to infer latent utilities (see Section 5.1). The elicitation method, query presentation, and context thus play a critical role in shaping measured preferences, compounded by cognitive biases and irrationalities.
The influence of psychological phenomena on preference learning has been well-documented in the literature, especially within the context of explicit preference elicitation for recommender systems. For instance, Tran et al. (2021) provide a thorough discussion of the relationship between psychology and recommender systems. Similarly, Atas et al. (2021) review how preference construction is influenced by cognitive biases, personality traits, and emotional states in recommender systems, discussing effects like serial position, framing, anchoring, choice overload, and preference visibility. In a more specialized discussion, Mandl et al. (2011) focus on cognitive biases in the context of consumer decision-making and its interaction with recommender systems. Mandl et al. (2011) specifically address cognitive biases in consumer decision-making in interaction with recommender systems. Finally, Kaufmann et al. (2023) link these psychological aspects to RLHF, discussing the common practice of using synthetic instead of real human feedback for algorithm evaluation and highlighting the limitations of that approach. They further discuss challenges posed by real human feedback, many of which are related to the concepts discussed in the following paragraphs, as well as the opportunities provided by integrating psychological insights into RLHF systems.
@@
Ordering effects pose another challenge in preference elicitation, where the sequence of queries can affect responses. Day et al. (2012) outline several factors contributing to these effects: institutional learning, changing preferences, and varying levels of cognitive effort. Institutional learning involves gaining familiarity with the task and feedback type, which can enhance labeler’s expertise and, consequently, the accuracy of their responses. However, due to the constructive nature of preferences, their preferences may evolve during the elicitation process, leading to changing preferences. This evolution might also be influenced by anchoring effects, where previously seen instances bias responses. Furthermore, cognitive effort levels can fluctuate due to factors like fatigue or boredom. This is closely related to choice overload, a form of fatigue from excessive choices, as discussed by Atas et al. (2021) and bounded rationality, as explored by Chen et al. (2013). In such scenarios, labelers might opt out of making a choice when overwhelmed by options. Bounded rationality refers to the limitations in human decision-making capabilities, particularly when processing large amounts of information, which aligns with the concept of choice overload. To address these challenges, studies like Bıyık et al. (2019) and Zhang et al. (2022) propose methods to reduce cognitive effort in responding to queries. Bıyık et al. (2019) focus on posing queries that are straightforward for humans to answer, while Zhang et al. (2022) enhance the human evaluation process by presenting queries in a user-friendly format.
@@
When multiple labelers collaborate on the same task in preference elicitation, as is studied, e.g., by Barnett et al. (2023) and Daniels-Koch & Freedman (2022), this may lead to another set of biases if they have the opportunity to exchange information. This exchange may either be direct or indirect through observing the system’s predictions, which are based on the other labeler’s feedback. Such interactions can affect their preferences through several mechanisms, as identified by Atas et al. (2021): anchoring effects, transfer of emotional states, and conflict avoidance. Anchoring effects, for instance, occur when a labeler’s choices are influenced by the knowledge of others’ preferences or system predictions, a phenomenon also discussed under the term preference visibility. This bias can lead labelers to align their preferences with the anchors they are exposed to, which is a significant consideration in recommender systems. Understanding these biases is crucial for designing RLHF systems that mitigate the influence of labeler interactions on preference construction.
@@
The effects previously discussed stem from systemic biases in preference expression. In addition to these biases, choices may also be affected by noise. This is commonly discussed under the term stochastic rationality, where an agent’s behavior is rational with respect to an unobserved random state. The reward-rational implicit choice framework, as introduced by Jeon et al. (2020), addresses this by integrating a rationality factor $\beta$ into the human choice model (see Eq. 3). This factor’s impact has been further examined by Ghosal et al. (2023) through synthetic experiments and user studies, demonstrating that accurately estimating this type-dependent rationality coefficient can enhance learning performance and guide feedback type selection (see Section 4.1.2). However, a practical method for estimating this factor remains a challenge. While Ghosal et al. (2023) use calibration feedback with a known latent utility function for estimation, such an approach is not feasible for most tasks. In a related study, Daniels-Koch & Freedman (2022) investigate a scenario with multiple teachers, focusing on the agent’s ability to select the most knowledgeable or rational teacher. Therefore, developing more advanced methods to estimate this factor, along with understanding its variability due to factors like fatigue or other ordering effects, presents a vital area for future research in preference elicitation

### v3

@@ 3.2.3 Representation-Specific Classes
While many approaches either rely on hard-coded features or learn a model entirely end-to-end, it is also possible to actively elicit new features from human feedback. Feature traces were proposed by Bobu et al. (2022) as an approach to actively learn new relevant features. This type of feedback relies on a human operator to demonstrate a behavior in which a certain feature of interest, such as the distance to a sensitive object, monotonically increases or decreases, e.g., by moving a robot arm closer to the object. They make it possible to extend the set of features once the current set can no longer adequately explain the human feedback supplied through another type of feedback. The defining characteristics of feature traces are that they are of descriptive (intent) and explicitly (explicitness) given in an active manner (involvement) for a single (arity) segment (granularity) on an instance-level abstraction.
Feature traces are strongly related to inter-temporal preferences (Section˜3.2.1) since both types rely on changes in feature or reward values in the course of a single trajectory. Bobu et al. (2022) propose to learn from feature traces by leveraging a Bradley-Terry model to learn the feature values, similar to other approaches that use such a model to learn reward values. Similar to importance feedback, feature traces rely on another type of feedback to actually make use of the learned features and is, therefore, a purely supplementary form of feedback. For instance, Bobu et al. (2022) use intervention feedback to train a reward model on the set of features derived using feature traces.
@@
Similarity queries are a feedback type aimed at learning a representation conforming to a notion of similarity and difference in the trajectory space. That aim is closely aligned with that of feature queries, though the actual queries are more similar to comparisons. The queries consist of triples of trajectories, with one anchor and two alternatives, for which the human has to decide which pair is more similar. Responses to similarity queries are given on observed behavior (involvement) with ternary arity, descriptive intent, and explicit feedback (explicitness), while the granularity and abstraction may vary. This type of feedback was first introduced by Bobu et al. (2023), who used it to learn representations for reward learning.
@@ 3.3 Initializations
Some modes of communicating reward functions are not interactive nor online and, therefore, do not directly fit within the scope of this survey (Section˜1.3). However, since these are often used to initialize a reward function for later refinement with some of the previously discussed interactive feedback types, they are still worth mentioning.
Initializations are most commonly given by examples of successful task completions, either in the form of terminal or goal states (Xie et al., 2018), expert demonstrations (Ibarz et al., 2018; Fu et al., 2018a; Palan et al., 2019; Lee et al., 2021b; Bıyık et al., 2022a; Abramson et al., 2022; Huang et al., 2023), demonstrations with success labels (Du et al., 2023), or ranked demonstrations (Brown et al., 2019). It is even possible to infer some human preferences by the state of the environment alone, e.g., assuming the parts of the initial environment that are under the user’s control are largely set up in accordance with the user’s preferences (Shah et al., 2019; Lindner et al., 2021a). Since offline initialization is not the main focus of our survey, we do not cover these works in detail. We refer the interested reader to literature on inverse RL (Arora & Doshi, 2021) for further details on learning from demonstrations in particular.
@@ 3.4 Choice of Feedback Type
The best feedback type is not always clear and may depend on the task, the user, or the agent. It may also change over time as the agent learns more about the task. Table˜2 may serve as a starting point to select a set of feedback types which may be applicable based on the possible user interaction and expertise, e.g., by the desired granularity, level of abstraction or involvement. This choice may be informed, among other features, by the task complexity or time horizon: Jain et al. (2015) compare purely passive (re-ranking based) with active (slight trajectory improvements) feedback and conclude that the former is usually preferred by humans when the action space is large while the latter is preferred for complex tasks. Similarly, Sumers et al. (2023) empirically demonstrate that language is a more effective teaching modality than demonstrations for complex concepts. The relevance of the time horizon is highlighted in the work by Sumers et al. (2022b), which suggests that instructive feedback is more efficient for short time horizons and descriptive feedback for longer horizons.
In addition to these static choices, Section˜4.1.2 will discuss how to choose a feedback type adaptively. It is also possible to let the user choose the feedback type, as demonstrated by Jain et al. (2015) who let the user choose between re-ranking and improvement. Jeon et al. (2020) propose to use this choice of feedback type itself as a source of information (“meta-choice”).
@@ 3.5 Combination of Feedback Types
The two-phase setup can be used to either initialize the reward model from offline data or to learn a representation that improves later reward learning. A common approach involves using one feedback type, typically demonstrations (see Section˜3.3), to initialize the reward model, subsequently fine-tuning this model with another type, such as comparisons. This method is exemplified by Ibarz et al. (2018), who combined demonstrations and pairwise comparisons. For a more detailed discussion on feedback types suitable for initialization, we refer to Section˜3.3. Alternatively, a representation-focused type of feedback might be employed initially (Section˜3.2.3) to cultivate a superior representation, followed by the application of a primary feedback type for reward model learning.
Combining a primary feedback type with a supplementary one can be beneficial to make the most of the available user interactions. Supplementary feedback, while not sufficient for reward model training by itself, can often be collected cheaply or as a side-effect of other interactions, making it a valuable addition to the primary feedback type. We refer to Section˜3.2.2 for a discussion on supplementary feedback types.
@@ 4 Label Collection
Figure 3: RLHF diagram highlighting components discussed in this section.
@@
Reddy et al. (2020) Ensemble-averaged KL-
@@
Hu et al. (2024b) Random
In this section, we explain how preference data can be collected for training a reward model, independent of the specific type of feedback used. We start by overviewing the active learning problem posed by query selection, e.g., how queries can be generated for a given query type and how even the type of feedback itself can be selected. Following this, we discuss issues arising in such human-computer interaction.
4.1 Active Learning
We first discuss query generation and selection for a given feedback type, then the extension of these techniques to the choice of feedback type.
4.1.1 Query Generation and Selection
As the RL agent is trained and its learned policy changes, the trajectories it generates will naturally evolve. Most work directly uses the trajectories obtained during RL training for preference learning. Using such trajectories, candidate queries are often generated randomly. However, for pairwise comparisons, a more efficient approach is to generate queries by exploiting preference transitivity (Hwang et al., 2023). Alternatively, the agent can also generate trajectories specifically to be used for querying (not necessarily for RL training), possibly with a learned transition model to limit the sampling cost (e.g., Reddy et al. (2020); Liu et al. (2023b)). This kind of active generation of queries can possibly lead to more informative ones.
In order to efficiently learn a suitable reward model, the agent must generate and select queries (Line 5 from Algorithm˜1) so that it can quickly learn a good strategy using those queries. This selection is performed via a criterion, usually called acquisition function [^2] We follow the terminology used in Bayesian active learning., which allows the queries to be compared. Although most existing work in RLHF uses an acquisition function that provides some measure of uncertainty (e.g., about the learned rewards), which is arguably one of the most important factors in active learning, an efficient acquisition function (see Table˜3 for works dedicated to improving query selection) may need to include various additional other aspects, such as: on-policy data, query simplicity, trajectory quality, query diversity, or query cost, which will be discussed one by one in the following. As a side note, interestingly, as highlighted by Habibian et al. (2022), the queries asked by an RL agent may also reveal its current reward learning stage.
@@ 4.1.1 Query Generation and Selection
Finally, it should be mentioned that naturally approaches with theoretical guarantees usually also use uncertainty as the main criterion for label collection (e.g., Novoseller et al. (2023), see Section˜7 for more details).
@@
Only focusing on uncertainty is likely insufficient or inefficient in RL because the previous methods may focus on choosing queries to identify the reward function as precisely as possible uniformly on the whole state-action space. However, it may be important to favor more on-policy trajectories to guarantee the relevance of the generated queries for the current policy, assuming that the behavior policy is stochastic (i.e., the current stochastic policy in SAC or the current deterministic policy with some noise, such as in DDPG). Indeed, improving reward learning in state-action regions that may never be visited with the current policy would lead to wasteful queries (Lindner et al., 2021b). One simple approach to ensure that the data is more on-policy is by simply sampling from the current policy (Cui & Niekum, 2018) or favoring more recently-generated trajectories (Hu et al., 2024b).
@@
When asking many queries (in batch, in sequence), the diversity of the queries becomes especially crucial to avoid asking redundant queries. Most work (Christiano et al., 2017; Lee et al., 2021b; Verma & Metcalf, 2024) follows a very myopic approach: Queries are often selected from a usually randomly-generated set of potential queries, and sequences of queries are not really coordinated. While some work exists that specifically tackles the selection of a batch of diverse queries (Bıyık & Sadigh, 2018; Bıyık et al., 2024), the latter is rarely considered due to its computational intractability. Indeed, planning ahead a sequence of queries would amount to solving a sequential decision-making problem under uncertainty over a combinatorial action space (i.e., the set of possible queries). For diverse batch querying, previous work considered using clustering methods such as k-medoids (Bıyık & Sadigh, 2018) or more recently determinantal point processes, which define probability distributions that promote diversity (Bıyık et al., 2024). In addition to approaches focused on selecting diverse queries, another promising research direction involves generating inherently diverse and informative behaviors to select from. This can be achieved with techniques such as entropy-regularized RL (Ziebart et al., 2008; Haarnoja et al., 2017) or generative methods based on flow networks (Bengio et al., 2021). The latter is particularly relevant for RLHF for generative models such as LLMs (Hu et al., 2024a). To our knowledge, this direction has not yet been explored in the context of RLHF.
@@
Since various different acquisition functions have been considered, some effort (Lee et al., 2021b; a) has been made to compare them. Generally speaking, uncertainty-based criteria (e.g., variance or average entropy) seem to often perform better empirically compared to random selection, a query diversity-based criterion alone or combined with an uncertainty-based criterion. Surprisingly, random selection has been shown to perform competitively in some cases (Christiano et al., 2017; Ibarz et al., 2018). Thus, a better understanding of which acquisition function should be preferred in which situation or domain is still an open question.
@@
where a higher $\beta$ corresponds to a more reliable human (see Section˜5.1.2). The setting in which this assumption does not hold, i.e., the labeler’s reward functions differ (a setting already considered in inverse RL (Choi & Kim, 2012)), has also been studied recently (Siththaranjan et al., 2024; Xue et al., 2024; Dong et al., 2024b; Myers et al., 2021; Bakker et al., 2022). Interestingly, a noisy oracle may sometimes provide more information than a completely reliable oracle. For instance, in Eq.˜3, the probability of erroneous answers given by the noisy oracle is related to how much a segment is preferred to the other one, and Chan et al. (2021) show that such Boltzmann-rational behavior can be more informative for reward inference than perfectly rational behavior. In contrast, only a binary preorder over segments can be inferred from the answers of a reliable and deterministic oracle, which may not be enough to recover the true reward function.
Another notable recent work is that by Ellis et al. (2024) who raise the issue of identifiability of the ground-truth reward function: many reward functions result in the same optimal behaviors. The authors propose a framework that enables the generation of acquisition function for various definitions of reward similarity, such as the one discussed in Section˜5.3.
@@ 4.1.2 Adaptive Choice of Feedback Type
$x=\operatornamewithlimits{argmax}_{x\in\mathcal{X}}\mathbb{E}_{c_{h}\sim P(c_{h}\mid x)}\big[D_{\mathit{KL}}\big(\mathbb{P}(\theta\mid c_{h},x)\mid\mathbb{P}(\theta)\big)\big]\,\text{,}$
@@
Ghosal et al. (2023) find that the most informative feedback type depends on the (type-dependent) rationality of the human labeler (see Section˜4.2.1). More precisely, it is shown that the most informative feedback depends on the rationality factor, e.g., while demonstrations are more informative than comparisons when the human is highly rational, comparisons should be preferred in less-rational settings. Given that this rationality might change due to factors such as fatigue or an individual labeler’s capabilities, this suggests that adaptively adjusting the feedback type during the labeling process may be worthwhile. Further study of this relationship is a promising area for future work.
@@ 4.2.1 Psychology-Aware Preference Elicitation
Preference learning methods typically assume the existence of inherent, stable preferences that can be elicited through querying. Contrary to this assumption, psychological research, such as the work by Lichtenstein & Slovic (2006), indicates that preferences are often constructed during the elicitation process and may vary with the method of elicitation or over time. This suggests that the feedback type not only affects elicitation’s effectiveness but also shapes preferences. Systematic biases, noise, and other psychological factors may influence observed choices, challenging the traditional models of human choice used to infer latent utilities (see Section˜5.1). The elicitation method, query presentation, and context thus play a critical role in shaping measured preferences, compounded by cognitive biases and irrationalities.
The influence of psychological phenomena on preference learning has been well-documented in the literature, especially within the context of explicit preference elicitation for recommender systems. For instance, Tran et al. (2021) provide a thorough discussion of the relationship between psychology and recommender systems. Similarly, Atas et al. (2021) review how preference construction is influenced by cognitive biases, personality traits, and emotional states in recommender systems, discussing effects like serial position, framing, anchoring, choice overload, and preference visibility. In a more specialized discussion, Mandl et al. (2011) focus on cognitive biases in the context of consumer decision-making and its interaction with recommender systems. Finally, Kaufmann et al. (2023) link these psychological aspects to RLHF, discussing the common practice of using synthetic instead of real human feedback for algorithm evaluation and highlighting the limitations of that approach. They further discuss challenges posed by real human feedback, many of which are related to the concepts discussed in the following paragraphs, as well as the opportunities provided by integrating psychological insights into RLHF systems.
@@
Ordering effects pose another challenge in preference elicitation, where the sequence of queries can affect responses. Day et al. (2012) outline several factors contributing to these effects: institutional learning, changing preferences, and varying levels of cognitive effort. Institutional learning involves gaining familiarity with the task and feedback type, which can enhance labelers’ expertise and, consequently, the accuracy of their responses. However, due to the constructive nature of preferences, these preferences may evolve during the elicitation process, leading to changing preferences. This evolution might also be influenced by anchoring effects, where previously seen instances bias responses. Furthermore, cognitive effort levels can fluctuate due to factors like fatigue or boredom. This is closely related to choice overload, a form of fatigue from excessive choices, as discussed by Atas et al. (2021) and bounded rationality, as explored by Chen et al. (2013). In such scenarios, labelers might opt out of making a choice when overwhelmed by options. Bounded rationality refers to the limitations in human decision-making capabilities, particularly when processing large amounts of information, which aligns with the concept of choice overload. To address these challenges, studies like Bıyık et al. (2019) and Zhang et al. (2022) propose methods to reduce cognitive effort in responding to queries. Bıyık et al. (2019) focus on posing queries that are straightforward for humans to answer, while Zhang et al. (2022) enhance the human evaluation process by presenting queries in a user-friendly format.
@@
Multiple labelers often collaborate on the same task in preference elicitation, as studied, e.g., by Barnett et al. (2023) and Daniels-Koch & Freedman (2022). This collaboration may lead to another set of biases if they have the opportunity to exchange information. The exchange may either be direct or indirect through observing the system’s predictions, which are based on the other labeler’s feedback. Such interactions can affect their preferences through several mechanisms, as identified by Atas et al. (2021): anchoring effects, transfer of emotional states, and conflict avoidance. Anchoring effects, for instance, occur when a labeler’s choices are influenced by the knowledge of others’ preferences or system predictions, a phenomenon also discussed under the term preference visibility. This bias can lead labelers to align their preferences with the anchors they are exposed to, which is a significant consideration in recommender systems. Understanding these biases is crucial for designing RLHF systems that mitigate the influence of labeler interactions on preference construction.
@@
The effects previously discussed stem from systemic biases in preference expression. In addition to these biases, choices may also be affected by noise. This is commonly discussed under the term stochastic rationality, where an agent’s behavior is rational with respect to an unobserved random state. The reward-rational implicit choice framework, as introduced by Jeon et al. (2020), addresses this by integrating a rationality factor $\beta$ into the human choice model (see Eq.˜3). This factor’s impact has been further examined by Ghosal et al. (2023) through synthetic experiments and user studies, demonstrating that accurately estimating this type-dependent rationality coefficient can enhance learning performance and guide feedback type selection (see Section˜4.1.2). However, a practical method for estimating this factor remains a challenge. While Ghosal et al. (2023) use calibration feedback with a known latent utility function for estimation, such an approach is not feasible for most tasks. In a related study, Daniels-Koch & Freedman (2022) investigate a scenario with multiple teachers, focusing on the agent’s ability to select the most knowledgeable or rational teacher. Therefore, developing more advanced methods to estimate this factor, along with understanding its variability due to factors like fatigue or other ordering effects, presents a vital area for future research in preference elicitation.